Find two unit vectors orthogonal to both ü = (2, -2, -3) and (0,3,6). unit vectors
Answer #1

Slution Given ü (2,-2,-3) = {0, 3,67 Required two unit vectors olthefend to both u and ris U XV + luxull U= (2-2, -37, <0, 3,- ) 1-4 I tik unit vectory = + dil -1-45+2k and unit vector = _i_49+2k 1 +49_2k V21 2 va